OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (–2, 2) and is parallel to y = 4x + 7.

OpenStudy (andriod09):

first thing is first, do you understand the slop intersecept concept?

OpenStudy (anonymous):

y = 4x + 10 y = 1/4 x + 5/2 y = -4x - 6 y = – 1/4 x + 3/2

OpenStudy (anonymous):

somewhat

OpenStudy (andriod09):

\[\frac{ \Delta y }{ \Delta x }\]

OpenStudy (andriod09):

that is the change of y, over the change of x

OpenStudy (andriod09):

so type the equation using \[\frac {\Delta y} {\Delta x}\]

OpenStudy (anonymous):

example please? /:

OpenStudy (andriod09):

lets take you last answer and try it. okay?

OpenStudy (anonymous):

okay

OpenStudy (andriod09):

\[\frac {\Delta y} {\Delta x} \] \[y=\frac {-1} {4}x + \frac {3} {2}\]

OpenStudy (anonymous):

2/6

OpenStudy (andriod09):

but first you have to find the y

OpenStudy (andriod09):

so find the why first.

OpenStudy (anonymous):

wouldn't it be 2? -1+3

OpenStudy (andriod09):

yup! that 's correct, now fill in \[\frac {\Delta y} {\Delta x}\]

OpenStudy (anonymous):

I still stick with my guess of 2/6

OpenStudy (anonymous):

I know that my answer isn't correct, but I'm not sure what I'm doing wrong

OpenStudy (andriod09):

no, you have to fill out the equation. use the equation button to fill it out

OpenStudy (anonymous):

rephrase please?

OpenStudy (anonymous):

im sorry i just dont get it

OpenStudy (andriod09):

okay, put the answer you got for y, in the \[\Delta y\] and then what you got for x in the \[\Delta x\]

OpenStudy (anonymous):

can you show me?

OpenStudy (andriod09):

yeah. give me your x and your y first and i'll show you.

OpenStudy (anonymous):

2/6

OpenStudy (andriod09):

is that your x or your y?

OpenStudy (anonymous):

i would say that 2 is my x and 6 is my y but im assuming that isn't correct

OpenStudy (andriod09):

how about we start on the actual problem, it would be alot less confusing.

OpenStudy (anonymous):

ok

OpenStudy (andriod09):

so show me the \[\frac {\Delta y} {\Delta x}\] with the shown intergers

OpenStudy (anonymous):

I'm sorry but I really don't understand any of this. Can you tell me the answer and explain why it's so?

OpenStudy (andriod09):

wow i just realized something, you just have to graph the equation and then plot the points. -.-

OpenStudy (anonymous):

yeahh lol i think we do too

OpenStudy (andriod09):

so graph them and tell me what you get. BTW: try answers b and c. :D

OpenStudy (anonymous):

C?

OpenStudy (andriod09):

:D correct.

OpenStudy (anonymous):

Thanks! Can you please help with one more?

OpenStudy (andriod09):

#2

OpenStudy (anonymous):

brb for 5 mins

OpenStudy (anonymous):

Thanks!

OpenStudy (andriod09):

NP!